# Frequency-Guided Masking for Enhanced Vision Self-Supervised Learning

### v1

@@ Abstract
We present a novel frequency-based Self-Supervised Learning (SSL) approach that significantly enhances its efficacy for pre-training. Prior work in this direction masks out pre-defined frequencies in the input image and employs a reconstruction loss to pre-train the model. While achieving promising results, such an implementation has two fundamental limitations as identified in our paper. First, using pre-defined frequencies overlooks the variability of image frequency responses. Second, pre-trained with frequency-filtered images, the resulting model needs relatively more data to adapt to naturally looking images during fine-tuning. To address these drawbacks, we propose FOurier transform compression with seLf-Knowledge distillation (FOLK), integrating two dedicated ideas. First, inspired by image compression, we adaptively select the masked-out frequencies based on image frequency responses, creating more suitable SSL tasks for pre-training. Second, we employ a two-branch framework empowered by knowledge distillation, enabling the model to take both the filtered and original images as input, largely reducing the burden of downstream tasks. Our experimental results demonstrate the effectiveness of FOLK in achieving competitive performance to many state-of-the-art SSL methods across various downstream tasks, including image classification, few-shot learning, and semantic segmentation.
@@ 1 Introduction
In recent years, Self-Supervised Learning (SSL) has gained considerable interest in the context of visual pre-training. This interest stems from its prominent capability of extracting meaningful visual representations from the vast expanse of readily available, unlabeled images without the need for costly manual labeling (Ben-Shaul et al., 2024; Su et al., 2024; Almalki and Latecki, 2024). Key to this advancement is several pre-training methods established with different pretext tasks, including multi-view contrastive learning (Oord et al., 2018; Chen et al., 2020b; Tian et al., 2020b; He et al., 2020), Masked Image Modeling (MIM) (Bao et al., 2022; He et al., 2022a; Xie et al., 2022; Monsefi et al., 2024a; Oquab et al., 2024), Masked Frequency Modeling (MFM) (Xie et al., 2023; Liu et al., 2023; Zheng et al., 2024), and self-supervised Knowledge Distillation (KD) (Kakogeorgiou et al., 2022; Zhou et al., 2022; Chen et al., 2020c; Chen and He, 2021; Caron et al., 2021). In the recent popular approach of Masked Image Modeling (MIM), a key strategy involves masking portions of an image and then tasking models with either reconstructing these hidden sections or generating feature representations for them (Bao et al., 2022; Xie et al., 2022; He et al., 2022a; Yi et al., 2023). Through this process, the model is encouraged to learn robust feature representations that capture the underlying structure between unmasked and masked image parts, thereby enhancing its understanding of image semantics.
Instead of masking in the spatial domain (MIM), Masked Frequency Modeling (MFM) (Xie et al., 2023) introduced a self-supervised approach that masks frequency components of the input image. Since high-level semantics and low-level details of an image can be separated into different frequency components (Oppenheim and Lim, 1981; Piotrowski and Campbell, 1982; Navard and Yilmaz, 2024), the frequency domain offers a more convenient avenue for revealing underlying image patterns. The pretext task in MFM is to predict the masked frequencies from the frequency-filtered image (see Section 3.1). It has two main advantages: first, it can help avoid issues encountered when analyzing raw pixel values in the spatial domain, such as spatial redundancy (Wang et al., 2020; Chen et al., 2023); second, unlike the patch-based masking used in MIM which often restricts the model to a Vision Transformer (ViT), MFM is suitable for both ViT and Convolutional Neural Network (CNN)-based models.
@@ 2.1 Self-supervised Learning
Self-supervised Learning (SSL) methods have been developed to exploit large-scale unlabeled data for learning discriminative representations, which can then benefit a variety of downstream tasks (Chong et al., 2023). Early SSL approaches rely on several pretext tasks, such as rotation prediction (Gidaris et al., 2018), jigsaw puzzle (Noroozi and Favaro, 2016), and colorization (Zhang et al., 2016). A branch of more recent studies follows a contrastive SSL paradigm (Chen et al., 2020b, c; Ci et al., 2022; Chen et al., 2020d; Tian et al., 2020b; Perera et al., 2024). SimCLR (Chen et al., 2020b, c; Monsefi et al., 2024b; Zhou et al., 2024) considers two augmentations of a given image as positives and the augmentations of all other images in the batch as negatives, on top of which a contrastive loss is utilized for model learning. MOCO (He et al., 2020; Chen et al., 2020d) maintains a dynamic dictionary of encoded representations with a momentum-updated encoder to generate consistent embeddings for contrastive learning. Instead of relying on negative samples and large training batches, BYOL (Grill et al., 2020) adopts a teacher-student framework that enforces consistency between representations of two augmented views of the same image generated by the two models. More recently, Correlational Image Modeling (CIM) (Li et al., 2023) operates by predicting correlation maps between randomly cropped image regions (exemplars) from a given image (context). It employs a bootstrap learning architecture with online and target encoders and a simple cross-attention mechanism to process the exemplars and context.
@@ 2.2 Masked Image Modeling
Masked Image Modeling (MIM) is an exciting approach to self-supervised visual learning. The central concept is to rebuild or reconstruct the hidden parts of images or to predict general characteristics like the image’s category (Bao et al., 2022; Zhou et al., 2022; Chen et al., 2020a; Xie et al., 2022; Wei et al., 2022). It draws inspiration from the success of masked language modeling in natural language processing (Devlin et al., 2018; Liu et al., 2019), adapting the concept to the visual domain. Specifically, BEiT (Bao et al., 2022) utilizes a discrete pre-trained Variational AutoEncoder (VAE) to create discrete tokens for image patches, then it tasks the model with predicting such discrete tokens of masked patches in images. The iBOT method (Zhou et al., 2022) offers improvements over BEiT by adopting a teacher-student self-distillation strategy where the teacher model simultaneously serves as an online tokenizer, instead of using a pre-trained discrete VAE. MAE (He et al., 2022a) takes a more aggressive masking strategy, with typically around 75% of patches being masked, and recovers the missing pixels using an autoencoder. To further investigate which parts of an image should be masked for efficient self-supervisory, AttMask (Kakogeorgiou et al., 2022) proposes the utilization of an attention map generated by a teacher model, and masks the highly attended patches from the image for the student model learning.
@@ 3.2.1 Informed Filters
The generation of and filters is illustrated in Fig. 2. The input image is first converted to grayscale to ensure the creation of a single common filter for all three RGB channels, as generating and applying filters separately to each channel can result in unnatural and corrupted visual information. The image is converted to a frequency spectrum using 2D FFT, and then the frequency components with the highest magnitudes are identified according to a threshold, which is uniformly sampled from a set of values ( in our experiments). The informed filters are then created to retain () or mask () these frequencies. The two filters are randomly selected with an equal possibility (i.e. ) and applied to the spectrum. By applying inverse FFT to the filtered spectrum, we restore a frequency-masked image which will serve as an input to the model during pre-training. It is important to note that and filters are uniquely generated based on individual images (examples are provided in Appendix C), which introduces greater variation in training samples and presents a harder training task compared to using constant filters. In addition to comparing our approach with the low/high-passed filters used in MFM (Xie et al., 2023), we also conducted an ablation study using a set of random frequency filters to demonstrate the effectiveness of our proposed informed filters (see Appendix B.3).
@@ 3.2.2 Making Backbone Familiar with Natural Images
To further improve the training efficiency and model robustness in masked frequency modeling, we incorporate a self-distillation design (Grill et al., 2020; Caron et al., 2021; Tarvainen and Valpola, 2017) in our proposed approach. The original MFM (Xie et al., 2023) method only tasked the model with the reconstruction of missing frequencies from the frequency-masked view of the image. Such an approach potentially overlooked the data distribution of the original image space, as the model only sees frequency-masked images in pre-training, resulting in higher data demands to adapt to naturally looking images during fine-tuning. To resolve this issue, we propose to properly inject the original image information into the training process via a self-distillation technique advocated by works such as BYOL (Grill et al., 2020) and DINO (Caron et al., 2021). Here, we present our FOLK framework, detailed in Fig. 3. Note that FOLK does not require additional training stages for pre-processing, such as the offline tokenizer employed by BEiT (Bao et al., 2022).
FOLK starts by generating two views, and , from an input image using distinct transformations to create varied perspectives. This follows the methodology employed by DINO (Caron et al., 2021) and the transformations include random cropping, color jittering, etc. Note that, unlike DINO or ATTMask (Kakogeorgiou et al., 2022), we do not utilize the concept of local views, which helps keep an efficient framework. After applying 2D FFT to the view (or ), the and filters are uniquely generated according to this view (detailed in Section 3.2.1 and Fig. 2). One filter is then randomly selected and applied to the frequency spectrum, and the retained frequency components are processed through inverse FFT to restore a frequency-masked view (or ), see Equation 2. The student model receives this frequency-masked view (or ) and predicts against two targets: reconstruction of the missing frequencies discarded by the filter, and reconstruction of the feature representation of the other original view (or ) generated by the teacher model.
@@
where . is the output dimension of , and is a temperature parameter. A similar formula holds for the teacher counterpart with and . The usage of a scaled softmax allows the sharpening of the output distribution (especially for the teacher) to avoid model collapse in practice (Caron et al., 2021). Additionally, we adhere to the centering methodology presented in DINO to further avoid model collapse and reduce large batch size dependency. Detailed descriptions of the heads are provided in Appendix A.3.
@@ 3.2.3 Comprehensive Loss Calculation
where a hyperparameter controls the weights between two loss terms, which is set as in our experiments, unless stated otherwise. Note that for a single input image, takes an average over two terms for the two different views. This comprehensive loss facilitates the simultaneous model learning on the two tasks introduced above, the masked frequencies reconstruction (through ) and original image feature reconstruction with self-distillation (through ). Furthermore, an ablation study for different choices of the hyperparameter is provided in Section 4.2.4.
@@ 4.1 Setup
In this study, we employ two well-established model architectures as the foundation for our experiments: the ViT-Small (ViT-S/16) (Dosovitskiy et al., 2021) and the ResNet-50 (He et al., 2016). These models are chosen for their proven effectiveness and versatility, showing the power of our model with different types of model architecture. We adopt the ImageNet-1K training dataset (Deng et al., 2009) without labels for pre-training our self-supervised learning approach.
Our model’s performance is evaluated across two critical areas: image classification and semantic segmentation. For image classification, we continue to leverage the ImageNet-1K dataset (Deng et al., 2009) to assess the generalizability and effectiveness of the learned features. In contrast, for semantic segmentation, we utilize the ADE20K dataset (Zhou et al., 2017), a standard benchmark in scene parsing and segmentation tasks. This bifurcated approach to evaluation ensures a thorough analysis of the models’ capabilities in varied contexts. Our computational infrastructure supports these extensive experiments, consisting of four nodes, each of which has four NVIDIA A100 80GB GPUs, in total 16 GPUs. Please see Appendix A for implementation details. Also, see Appendix B.4 for GPU usage.
@@ 4.2.1 Image Classification
In this section, we focus on the fine-tuning capabilities of different vision pre-training techniques, using the ViT-S/16 encoder on the ImageNet-1K dataset. The motivation behind this analysis stems from the need to understand how different SSL strategies, which range from traditional methods to novel approaches like the FOLK method introduced here, perform under uniform testing conditions with the well-established ImageNet benchmark dataset (Deng et al., 2009).
| Method | Ref | Data | Epoch | Token | ViT-S |
| Scratch | (Touvron et al., 2021) | - | - | - | 79.9 |
| MAE | (He et al., 2022a) | ImageNet-1K | 300 | ✓ | 80.6 |
| SimMIM | (Xie et al., 2022) | ImageNet-1K | 300 | ✓ | 80.9 |
| iBOT | (Zhou et al., 2022) | ImageNet-1K | 300 | ✓ | 81.1 |
| BEiT | (Bao et al., 2022) | ImageNet-1K | 300 | ✓ | 81.3 |
| + DALL-E | |||||
| AttMask | (Kakogeorgiou et al., 2022) | ImageNet-1K | 300 | ✓ | 81.3 |
| MoCo V3 | (Ci et al., 2022) | ImageNet-1K | 600 | - | 81.4 |
| DINO | (Caron et al., 2021) | ImageNet-1K | 1600 | - | 81.5 |
@@
| FOLK | Ours | ImageNet-1K | 800 | - | 82.1 |
In Table 1, notably, the traditional training from scratch method achieves a baseline accuracy of , which sets the stage for evaluating the added value of SSL strategies. Models employing masked token (i.e. image patch) strategies, such as MAE, SimMIM, BEiT, and AttMask, show a notable improvement over the baseline, with accuracies ranging from to . This highlights the potential benefits of using masked tokens in training to boost model performance. However, these methods only work with image token-based models, such as ViTs.
In contrast, the FOLK method demonstrates a compelling advancement in vision SSL without the need for token masking or additional datasets, like DALL-E (Ramesh et al., 2021) used in the BEiT model. The standard FOLK approach reaches an accuracy of with epochs pre-training, matching the highest performance of other models trained under similar conditions but with more epochs, i.e. MoCo V3 and DINO with and epochs, respectively. Its performance is further improved to when extended to 800 epochs pre-training, surpassing all other methods with a notable margin. This showcases FOLK’s superior capability, suggesting that its method of integrating learning from dual inputs—filtered and original images—significantly enhances learning efficiency and model efficacy. Furthermore, we provide a qualitative analysis with visualizations of our proposed filters on example images, along with the model predictions for the pretext task, which further demonstrate a successful pre-training achieved by FOLK. This is detailed in Appendix C.

### v2

@@ Abstract
We present a novel frequency-based Self-Supervised Learning (SSL) approach that significantly enhances its efficacy for pre-training. Prior work in this direction masks out pre-defined frequencies in the input image and employs a reconstruction loss to pre-train the model. While achieving promising results, such an implementation has two fundamental limitations as identified in our paper. First, using pre-defined frequencies overlooks the variability of image frequency responses. Second, pre-trained with frequency-filtered images, the resulting model needs relatively more data to adapt to naturally looking images during fine-tuning. To address these drawbacks, we propose FOurier transform compression with seLf-Knowledge distillation (FOLK), integrating two dedicated ideas. First, inspired by image compression, we adaptively select the masked-out frequencies based on image frequency responses, creating more suitable SSL tasks for pre-training. Second, we employ a two-branch framework empowered by knowledge distillation, enabling the model to take both the filtered and original images as input, largely reducing the burden of downstream tasks. Our experimental results demonstrate the effectiveness of FOLK in achieving competitive performance to many state-of-the-art SSL methods across various downstream tasks, including image classification, few-shot learning, and semantic segmentation. https://github.com/aminK8/FOLK
@@ 1 Introduction
In recent years, Self-Supervised Learning (SSL) has gained considerable interest in the context of visual pre-training. This interest stems from its prominent capability of extracting meaningful visual representations from the vast expanse of readily available, unlabeled images without the need for costly manual labeling (Ben-Shaul et al., 2024; Su et al., 2024; Almalki & Latecki, 2024). Key to this advancement is several pre-training methods established with different pretext tasks, including multi-view contrastive learning (Oord et al., 2018; Chen et al., 2020b; Tian et al., 2020b; He et al., 2020), Masked Image Modeling (MIM) (Bao et al., 2021; He et al., 2022a; Xie et al., 2022; Monsefi et al., 2024a; Oquab et al., 2024), Masked Frequency Modeling (MFM) (Xie et al., 2023; Liu et al., 2023; Zheng et al., 2024), and self-supervised Knowledge Distillation (KD) (Kakogeorgiou et al., 2022; Zhou et al., 2022; Chen et al., 2020c; Chen & He, 2021; Caron et al., 2021). In the recent popular approach of Masked Image Modeling (MIM), a key strategy involves masking portions of an image and then tasking models with either reconstructing these hidden sections or generating feature representations for them (Bao et al., 2021; Xie et al., 2022; He et al., 2022a; Yi et al., 2023). Through this process, the model is encouraged to learn robust feature representations that capture the underlying structure between unmasked and masked image parts, thereby enhancing its understanding of image semantics.
Instead of masking in the spatial domain (MIM), Masked Frequency Modeling (MFM) (Xie et al., 2023) introduced a self-supervised approach that masks frequency components of the input image. Since high-level semantics and low-level details of an image can be separated into different frequency components (Oppenheim & Lim, 1981; Piotrowski & Campbell, 1982; Navard & Yilmaz, 2024), the frequency domain offers a more convenient avenue for revealing underlying image patterns. The pretext task in MFM is to predict the masked frequencies from the frequency-filtered image (see Section 3.1). It has two main advantages: first, it can help avoid issues encountered when analyzing raw pixel values in the spatial domain, such as spatial redundancy (Wang et al., 2020; Chen et al., 2023); second, unlike the patch-based masking used in MIM which often restricts the model to a Vision Transformer (ViT), MFM is suitable for both ViT and Convolutional Neural Network (CNN)-based models.
@@ 2.1 Self-supervised Learning
Self-supervised Learning (SSL) methods have been developed to exploit large-scale unlabeled data for learning discriminative representations, which can then benefit a variety of downstream tasks (Chong et al., 2023). Early SSL approaches rely on several pretext tasks, such as rotation prediction (Gidaris et al., 2018), jigsaw puzzle (Noroozi & Favaro, 2016), and colorization (Zhang et al., 2016). A branch of more recent studies follows a contrastive SSL paradigm (Chen et al., 2020b; c; Ci et al., 2022; Chen et al., 2020d; Tian et al., 2020b; Perera et al., 2024). SimCLR (Chen et al., 2020b; c; Monsefi et al., 2024b; Zhou et al., 2024) considers two augmentations of a given image as positives and the augmentations of all other images in the batch as negatives, on top of which a contrastive loss is utilized for model learning. MOCO (He et al., 2020; Chen et al., 2020d) maintains a dynamic dictionary of encoded representations with a momentum-updated encoder to generate consistent embeddings for contrastive learning. Instead of relying on negative samples and large training batches, BYOL (Grill et al., 2020) adopts a teacher-student framework that enforces consistency between representations of two augmented views of the same image generated by the two models. More recently, Correlational Image Modeling (CIM) (Li et al., 2023) operates by predicting correlation maps between randomly cropped image regions (exemplars) from a given image (context). It employs a bootstrap learning architecture with online and target encoders and a simple cross-attention mechanism to process the exemplars and context.
@@ 2.2 Masked Image Modeling
Masked Image Modeling (MIM) is an exciting approach to self-supervised visual learning. The central concept is to rebuild or reconstruct the hidden parts of images or to predict general characteristics like the image’s category (Bao et al., 2021; Zhou et al., 2022; Chen et al., 2020a; Xie et al., 2022; Wei et al., 2022). It draws inspiration from the success of masked language modeling in natural language processing (Devlin et al., 2018; Liu et al., 2019), adapting the concept to the visual domain. Specifically, BEiT (Bao et al., 2021) utilizes a discrete pre-trained Variational AutoEncoder (VAE) to create discrete tokens for image patches, then it tasks the model with predicting such discrete tokens of masked patches in images. The iBOT method (Zhou et al., 2022) offers improvements over BEiT by adopting a teacher-student self-distillation strategy where the teacher model simultaneously serves as an online tokenizer, instead of using a pre-trained discrete VAE. MAE (He et al., 2022a) takes a more aggressive masking strategy, with typically around 75% of patches being masked, and recovers the missing pixels using an autoencoder. To further investigate which parts of an image should be masked for efficient self-supervisory, AttMask (Kakogeorgiou et al., 2022) proposes the utilization of an attention map generated by a teacher model, and masks the highly attended patches from the image for the student model learning.
@@ 3.2.1 Informed Filters
The generation of and filters is illustrated in Fig. 2, with the pseudocode available in Appendix A.4. The input image is first converted to grayscale to ensure the creation of a single common filter for all three RGB channels, as generating and applying filters separately to each channel can result in unnatural and corrupted visual information. The image is converted to a frequency spectrum using 2D FFT, after which the frequency components with the highest magnitudes are identified based on a threshold uniformly sampled from a predefined set of values ( in our experiments, detailed in Appendix Section B.5.1). The informed filters are then created to retain () or mask () these frequencies. The two filters are randomly selected with an equal possibility (i.e. ) and applied to the spectrum (effects of different selection probabilities are provided in Appendix Section B.5.2). By applying inverse FFT to the filtered spectrum, we restore a frequency-masked image which will serve as an input to the model during pre-training. It is important to note that and filters are uniquely generated based on individual images (examples are provided in Appendix D), which introduces more significant variation in training samples and presents a more challenging training task compared to using constant filters. In addition to comparing our approach with the low/high-passed filters used in MFM (Xie et al., 2023), we also conducted an ablation study using a set of random frequency filters to demonstrate the effectiveness of our proposed informed filters (refer to Appendix B.5.4).
@@ 3.2.2 Making Backbone Familiar with Natural Images
To further improve the training efficiency and model robustness in masked frequency modeling, we incorporate a self-distillation design (Grill et al., 2020; Caron et al., 2021; Tarvainen & Valpola, 2017) in our proposed approach. The original MFM (Xie et al., 2023) method only tasked the model with the reconstruction of missing frequencies from the frequency-masked view of the image. Such an approach potentially overlooked the data distribution of the original image space, as the model only sees frequency-masked images in pre-training, resulting in higher data demands to adapt to naturally looking images during fine-tuning. To resolve this issue, we propose to properly inject the original image information into the training process via a self-distillation technique (see Section 2.3 for more information about self-distillation). Here, we present our FOLK framework, detailed in Fig. 3. Note that FOLK does not require additional training stages for pre-processing, such as the offline tokenizer employed by BEiT (Bao et al., 2021).
FOLK starts by generating two views, , and , from an input image , following the same data augmentations adopted by DINO (Caron et al., 2021). After applying 2D FFT to the view (or ), the and filters are uniquely generated according to this view (detailed in Section 3.2.1 and Fig. 2). One filter is then randomly selected and applied to the frequency spectrum, and the retained frequency components are processed through inverse FFT to restore a frequency-masked view (or ), see Equation 2. The student model receives this frequency-masked view (or ) and predicts against two targets: reconstruction of the missing frequencies discarded by the filter, and reconstruction of the feature representation of the other original view (or ) generated by the teacher model.
@@
where . is the output dimension of , and is a temperature parameter. A similar formula holds for the teacher counterpart with and . To sharpen the output distributions—particularly for the teacher—we employ a scaled softmax function alongside the centering technique introduced in Caron et al. (2021). This combination helps prevent model collapse and reduces the dependency on large batch sizes during training. Detailed descriptions of the heads are provided in Appendix A.3.
@@ 3.2.3 Comprehensive Loss Calculation
where a hyperparameter controls the weights between two loss terms, which is set as in our experiments, unless stated otherwise. Note that for a single input image, takes an average over two terms for the two different views. This comprehensive loss facilitates the simultaneous model learning on the two tasks introduced above, the masked frequencies reconstruction (through ) and original image feature reconstruction with self-distillation (through ). Furthermore, an ablation study for different choices of the hyperparameter is provided in Section B.5.3.
3.2.4 Comparisons with Other Self-distillation-based Methods
Though FOLK deliberately leverages self-distillation to overcome the limitations of MFM — particularly its inability to expose the model to natural images — it is also important to clarify FOLK’s distinctions from other self-distillation-based methods such as DINO (Caron et al., 2021), AttMask (Kakogeorgiou et al., 2022), and iBOT (Zhou et al., 2022). In contrast to the spatial domain augmentation paradigms adopted by these methods (e.g. random cropping and masking on the original images), FOLK performs augmentation in the frequency domain by masking with adaptive filters and presenting frequency-masked images to the student model to enhance its visual understandings. Note that the input to the teacher model is only spatially augmented (i.e. random cropping), thereby preserving the model’s perception of natural images during pretraining. Moreover, FOLK incorporates a combination of both reconstructive and comparative losses, presenting rational yet more challenging pretraining tasks compared to approaches that rely solely on comparative losses, such as DINO and iBOT. Hence, we do not consider FOLK a close replication of previously presented methods. Please refer to Appendix Section C for a more detailed comparison.
@@ 4 Experiments
In this section, we detail the experimental setup and evaluate our proposed FOLK framework on classification tasks using both full fine-tuning and few-shot learning approaches. Additional experiments, including semantic segmentation as a downstream task and ablation studies, are provided in Appendix B for further insights and comprehensive results.
@@ 4.1 Setup
In this study, we employ three well-established model architectures as the foundation for our experiments: the ViT-Small (ViT-S/16), ViT-Base (ViT-B/16)(Dosovitskiy et al., 2021) and the ResNet-50 (He et al., 2016) (For ResNet-50 results see Appendix Section B.2). These models are chosen for their proven effectiveness and versatility, showing the power of our model with different types of model architecture. We adopt the ImageNet-1K training dataset (Deng et al., 2009) without labels for pre-training our self-supervised learning.
Our model’s performance is evaluated across two critical areas: image classification and semantic segmentation. For image classification, we continue to leverage the ImageNet-1K dataset (Deng et al., 2009) to assess the generalizability and effectiveness of the learned features. In contrast, for semantic segmentation, we utilize the ADE20K dataset (Zhou et al., 2017), a standard benchmark in scene parsing and segmentation tasks. This bifurcated approach to evaluation ensures a thorough analysis of the models’ capabilities in varied contexts. Our computational infrastructure supports these extensive experiments, consisting of four nodes, each of which has four NVIDIA A100 80GB GPUs, in total 16 GPUs. Please see Appendix A for implementation details.
@@ 4.2.1 Image Classification
In this section, we focus on the fine-tuning capabilities of different vision pre-training techniques using the ViT-S/16 and ViT-B/16 encoders on the ImageNet-1K dataset. The motivation behind this analysis stems from the need to understand how different SSL strategies, which range from traditional methods to novel approaches like the FOLK method introduced here, perform under uniform testing conditions with the well-established ImageNet benchmark dataset (Deng et al., 2009).
| Method | Ref | Data | Epoch | Token | ViT-S | ViT-B |
| Scratch | (Touvron et al., 2021) | - | - | - | 79.9 | 81.8 |
| MAE | (He et al., 2022a) | INet-1K | 300 | ✓ | 80.6 | 82.9 |
| iBOT | (Zhou et al., 2022) | INet-1K | 1600 | ✓ | 81.1† | 84.0‡ |
| BEiT | (Bao et al., 2021) | INet-1K | 300 | ✓ | 81.3 | 82.9 |
| + DALL-E | ||||||
| AttMask | (Kakogeorgiou et al., 2022) | INet-1K | 300 | ✓ | 81.3 | N/A |
| MoCo V3 | (Ci et al., 2022) | INet-1K | 600 | - | 81.4 | 83.2 |
| DINO | (Caron et al., 2021) | INet-1K | 1600 | - | 81.5 | 82.8 |
@@
| FOLK | Ours | INet-1K | 800 | - | 82.1 | 84.0 |
Table 1 presents the top-1 accuracy results of various SSL approaches fine-tuned on ImageNet-1K using both ViT-S/16 and ViT-B/16 encoders. Methods utilizing masked token strategies, such as MAE, BEiT, iBOT, and AttMask, consistently outperform the baseline, achieving up to with ViT-S and with ViT-B with huge number of pre-training epochs. These improvements highlight the effectiveness of masked token approaches in enhancing model performance, particularly when extensive pre-training epochs or additional data (e.g., DALL-E (Ramesh et al., 2021) for BEiT) are employed.
However, our proposed FOLK method demonstrates superior performance without relying on tokens or external datasets. With only 300 epochs of pre-training, FOLK achieves accuracy with ViT-S and with ViT-B, matching or, surpassing other methods that require more epochs or additional data. Extending the pre-training to 800 epochs, FOLK further improves the ViT-S and ViT-B accuracy to and , surpassing all other methods in the ViT-S and ViT-B category under similar conditions or lower number of pre-training epochs. These results showcase FOLK’s ability to enhance learning efficiency and model efficacy by integrating learning from dual inputs—filtered and original images—without the complexities of token masking or the need for external data sources.